OpenStudy (anonymous):

Find the exact value of the six trig functions of the given angle. 19pi/6

OpenStudy (anonymous):

sin=-1/2 cos=-(sqrt3)/2 tan=(sqrt3)/3 cot=sqrt3 sec=-(2sqrt3)/3 csc=-2

OpenStudy (anonymous):

correct?